OpenStudy (anonymous):

Subtract. 65.32 – 12.2

OpenStudy (anonymous):

65-12=53 0.32-2=0.12 So 53+0.12=...

OpenStudy (anonymous):

you just gotta add 53+0.12 and you've got your result